OpenStudy (anonymous):

okay so i was gone this day and have no clue what to do or anything so please help again....Determine a and k so the given points are on the graph of the function (1,7)(2,-23) ; y=a(x+1)^2+k

OpenStudy (anonymous):

ok here is the deal, we go slow. replace x by 1, set the result = 7 and write \[a(1+1)^2+k=7\] \[2a+k=7\]

OpenStudy (anonymous):

omg satellit73 i feel stupid thanks so much for all your help seriously.

OpenStudy (anonymous):

then repeat with x = 2 and result = -23 to get \[a(2+1)^2+k=-23\] \[9a+k=-23\] so far so good?

OpenStudy (anonymous):

no problem, we take it step by step

OpenStudy (anonymous):

yes this is helping keep going please

OpenStudy (anonymous):

so now we have two equations involving a and k \[2a+k=7\] \[9a+k=-23\] and we want "a" and "k"

OpenStudy (anonymous):

except i made a mistake, because \[1+1=2\] and \[2^2=4\] so the first one should be \[a(1+1)^2+k=7\] \[4a+k=7\] my fault sorry about that.

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

so we still have two equations \[9a+k=-23\] \[4a+k=7\] and so we can subtract the second one from the first to get one equation with one variable

OpenStudy (anonymous):

we will get by subtraction \[5a=-30\] and therefore \[a=-6\]

OpenStudy (anonymous):

now since we know \[a=-6\] we can substitute into either equation and solve for k. we can write \[4\times (-6)+k=7\] \[-24+k=7\] \[k=31\] i think that should work

OpenStudy (anonymous):

so basically the bigger of the two will go on top?

OpenStudy (anonymous):

it makes no difference really which one is larger. by larger i assume you mean the coefficient of the "a" term. we could have written \[4a+k=7\] \[9a+k=-23\] and subtracted to get \[-5a=30\] and still \[a=-6\] i just like working with positive coefficients to make life a tiny bit simpler

OpenStudy (anonymous):

i dont see where k=31 where the 31 comes from?

OpenStudy (anonymous):

ok lets go slow again. we agree that \[a=-6\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and we also know that \[4a+k=7\] yes?

OpenStudy (anonymous):

si

OpenStudy (anonymous):

and since \[a=-6\] this tells us that \[4\times (-6)+k=7\] by substituting -6 for "a" so we know \[-24+k=7\] and we can solve for k by adding 24 to both sides to get \[k=7+24\] \[k=31\]

OpenStudy (anonymous):

ahhhhh haaa okay i get it haha sorry

OpenStudy (anonymous):

no need for sorry, this is confusing because you have to remember how to solve a system of equations to do this seemingly unrelated problem

OpenStudy (anonymous):

THANK YOU SOOOOO MUCH IM SORRY IT SUCKS BEING GONE AND HAVING A TEST DO AT MIDNIGHT THAT HE DIDNT TELL ME ABOUT.

OpenStudy (anonymous):

so your "final answer" is \[y=-6(x+1)^2+31\] ad we can check that this is right if you like

OpenStudy (anonymous):

THANK YOU THANK YOU THANK YOU HAHA ITS RIGHT :):):)

OpenStudy (anonymous):

how do you have a test at midnight?

OpenStudy (anonymous):

on line school for night owls?

OpenStudy (anonymous):

DUE AT MIDNIGHT HE GAVE US A TAKE HOME TEST ONLINE

OpenStudy (anonymous):

nice. any more feel free to post

OpenStudy (anonymous):

I GO TO A HIGH SCHOOL THAT GAVE ALL STUDENTS THEIR OWN LABTOPS AND SO OUR CLASS ASSIGNMENTS AND TEST AND QUIZES ARE ALL ONLINE BUT HE TEACHES US OFFLINE. HE'S NOT A VERY GOOD TEACHER....THIS ISNT ACTUALLY KAITLYN EITHER ITS HER FRIEND BUT OCCASIONALLY SHE LETS ME USE THIS!